What is the easiest way to get Y without plugging in guesses.
Example
4^y="radical"2
I know y= 1/4 but whats the correct way of getting this answer?

Answer by jim_thompson5910(35256)

4^y = sqrt(2)

4^y = 2^(1/2)

(2^2)^y = 2^(1/2)

2^(2*y) = 2^(1/2)

2*y = 1/2

y = (1/2)*(1/2)

y = 1/4